### v2

# Counting fixed-point-free Cayley permutations

Abstract
Two-sort species yield differential equations for functional digraphs of Cayley permutations. From these we obtain an explicit formula for fixed-point-free Cayley permutations and conjecture that their proportion tends to $1/e$, as for permutations and endofunctions. Our approach also yields counting formulas when the functional digraph is a tree, forest, or connected.
1 Introduction
The hat-check problem lies at the intersection of combinatorics and probability theory. Pierre Rémond de Montmort [14] first described it in 1708 in his “Essai d’analyse sur les jeux de hasard”. The problem asks for the probability that a uniformly random permutation is a derangement, that is, fixed-point-free. Montmort and Bernoulli later provided solutions featuring a striking occurrence of Napier’s constant in combinatorics: The number of derangements of size $n$ is $n!\bigl(1/0!-1/1!+1/2!-\cdots+(-1)^{n}/n!\bigr)$ and hence, as $n\to\infty$, the probability converges to $1/e$.
A lesser-known yet easier-to-prove result is that endofunctions have the same asymptotic behavior. Here endofunctions are maps from $n]=\{1,2,\ldots,n\}$ to $[n]$. Among all $n^{n}$ endofunctions, $(n-1)^{n}$ have no fixed points. As $n$ approaches infinity, the quotient converges again to $1/e$. Indeed, $(n-1)^{n}/n^{n}=(1-1/n)^{n}\to 1/e$.
Cayley permutations lie between permutations and endofunctions and present a more challenging version of this problem. They are endofunctions whose image contains every integer from one to its maximum, or equivalently, they are the lexicographically smallest representatives of positive integer words modulo order isomorphism. Cayley permutations play a key role in recent extensions of permutation pattern theory [1, 6, 7, 8, 9].
The difficulty of Cayley-derangements—Cayley permutations with no fixed points—lies in their biased nature. In permutations and endofunctions, fixed points are uniformly distributed. In Cayley permutations, smaller values appear more often and are thus more likely to be fixed. Furthermore, the combinatorial description is less transparent: a permutation decomposes as its fixed points together with a derangement of the rest, leading directly to enumeration by subfactorials, but no such simple decomposition exists for Cayley permutations.
This paper provides a counting formula for Cayley-derangements (Theorem 5.6). Namely, the number of Cayley-derangements on $[n]$ is
$\sum_{r=0}^{n}\frac{|\mathrm{Der}[r]|}{r!}\sum_{i=0}^{n}i!\genfrac{\{}{\}}{}{}{{n}}{{i}}_{{r}}$
[…]
We use two-sort species to study functional digraphs of Cayley-derangements, deriving differential equations for the broader class of $R$-recurrent Cayley permutations. These include Cayley permutations whose functional digraph is a tree, a forest, or connected. The appearance of species and functional digraphs here echoes Joyal’s [13] proof of Cayley’s formula. We outline the paper below. Section 2 reviews combinatorial species and functional digraphs. Section 3 introduces Cayley permutations. Section 4 constructs $R$-recurrent functional digraphs, and Section 5 derives the differential equations yielding our counting formula. Section 6 extends Joyal’s bijection to the two-sort setting, while Section 7 provides complementary unisort equations. Section 8 studies statistics such as internal nodes and connected components. Section 9 discusses generalizations.
2 Combinatorial species and functional digraphs
A $\mathbb{B}$-species (or simply species) is a functor $F:\mathbb{B}\to\mathbb{B}$, where $\mathbb{B}$ is the category of finite sets with bijections as morphisms. A species defines both a class of labeled combinatorial structures—F[U]F[U]—and how relabeling of the underlying set affects these structures—the transport of structure $F[\sigma]$. Since the transport of structure is a bijection, the cardinality of $F[U]$ depends solely on the cardinality of $U$, not on the nature of the elements of $U$. The generating series of a species $F$ is the formal power series
$F(x)=\sum_{n\geq 0}|F[n]|\frac{x^{n}}{n!},$
where $F[n]$ is short for $F[[n]]$ and $n]=\{1,2,\dots,n\}$ (in particular, $0]=\emptyset$).
Throughout this paper we use the following species, whose formal definitions can be found in Joyal’s seminal paper [13] or in the book by Bergeron, Labelle and Leroux [2]. These are the main references for combinatorial species theory. Shorter introductions can be found in our papers [9, 10, 11].
$1=E_{0}$
is the species characteristic of the empty set.
$X=E_{1}$
is the species characteristic of singletons.
$E$
is the species of sets.
$L$
is the species of linear orders.
$\mathcal{S}$
is the species of permutations.
$\mathcal{C}$
is the species of cyclic permutations (cycles).
Their generating series are $1(x)=1$, $X(x)=x$,
$E(x)=e^{x},\;L(x)=\mathcal{S}(x)=1/(1-x)\,\text{ and }\,\mathcal{C}(x)=-\log\bigl(1-x\bigr).$
Two species $F$ and $G$ are (combinatorially) equal, written $F=G$, if there is a natural isomorphism between them. Equality implies $F(x)=G(x)$, but not conversely: $L(x)=\mathcal{S}(x)$ while $L\neq\mathcal{S}$.
The species $G$ is a subspecies of $F$, written $G\subseteq F$, if for each pair of finite sets $U,V$ and bijection $\sigma:U\to V$ we have $G[U]\subseteq F[U]$ and $G[\sigma]=F[\sigma]|_{G[U]}$. We denote by $F_{+}$ and $F_{n}$ the subspecies of $F$ consisting of $F$-structures on nonempty sets and on sets with cardinality $n$, respectively.
Species can be viewed as a categorification of exponential generating series: each operation on formal power series—sum, product, composition, differentiation—has a canonical lift to species that recovers the original operation upon taking generating series. We now describe these lifts, defining only the set of structures; the transport of structure is implicit in each case, and a formal treatment can be found in the book by Bergeron et al. [2].
Given two species $F$ and $G$, an $(F+G)$-structure is either an $F$-structure or a $G$-structure:
$(F+G)[U]=F[U]\sqcup G[U],$
where $\sqcup$ denotes disjoint union. An $(F\cdot G)$-structure on $U$ is a pair $(s,t)$ such that $s$ is an $F$-structure on a subset $U_{1}\subseteq U$ and $t$ is a $G$-structure on $U_{2}=U\setminus U_{1}$:
$(F\cdot G)[U]=\bigsqcup_{(U_{1},U_{2})}F[U_{1}]\times G[U_{2}],$
where $U=U_{1}\sqcup U_{2}$ and $\times$” denotes cartesian product.
An $(F\circ G)$-structure, where $G[\emptyset]=\emptyset$, is a generalized partition where each block carries a $G$-structure and the set of blocks is structured by $F$:
$(F\circ G)[U]=\bigsqcup_{\beta\in\mathrm{Par}[U]}F[\beta]\times\prod_{B\in\beta}G[B],$
where $\mathrm{Par}[U]$ denotes the set of partitions of $U$. For instance, the species of ballots (ordered set partitions) is defined by $\mathrm{Bal}=L\circ E_{+}$ and hence $\mathrm{Bal}(x)=1/(2-e^{x})$; here $|\mathrm{Bal}[n]|$ is the $n$th Fubini number (A000670 [15]).
The derivative $F^{\prime}$, defined by $F^{\prime}[U]=F[U\sqcup\{\star\}]$, adds an element to the underlying set. Pointing $F^{\bullet}$, defined by $F^{\bullet}[U]=F[U]\times U$, distinguishes an element of the set. The two are related by $F^{\bullet}=X\cdot F^{\prime}$.
The species of endofunctions is defined by
$\mathrm{End}[U]=\{f:U\to U\}\quad\text{and}\quad\mathrm{End}[\sigma](f)=\sigma\circ f\circ\sigma^{-1}$
for any finite sets $U$ and $V$ and any bijection $\sigma:U\to V$. Throughout this paper we identify an endofunction $f\in\mathrm{End}[U]$ with its functional digraph, defined as the directed graph with vertex set $U$ and directed edges $(u,f(u))$, or $u\mapsto f(u)$, $u\in U$. Alternatively, the species of functional digraphs is the subspecies of directed graphs where each node has outdegree one. The transport of structures is inherited from the transport on directed graphs: a directed graph on $U$ with edges $\{(x,y)\}$ is mapped by the transport of structure along $\sigma$ to the directed graph on $V$ with edges $\{(\sigma(x),\sigma(y)\}$. Identifying an endofunction with its functional digraph is a natural isomorphism, and thus the species of endofunctions and functional digraphs are combinatorially equal. An instance of a functional digraph can be found in Figure 1.
337755441010141411111313669911121222881515
Figure 1: The functional digraph of $f=985776326459548\in\mathrm{End}[15]$, where the $i$th letter of $f$ is $f(i)$. Internal nodes are black while leaves are white.
Let $u\in U$ be a node in the functional digraph of $f$. We say that $u$ is internal if it has positive indegree, that is, if $u\in\mathrm{Im}(f)$. Otherwise, $u$ is a leaf. A fixed point of $f$ is an element $u\in U$ such that $f(u)=u$, which corresponds to a loop in the functional digraph.
The species of permutations $\mathcal{S}\subseteq\mathrm{End}$ consists of bijective endofunctions, and cycles $\mathcal{C}\subseteq\mathcal{S}$ are permutations whose functional digraph is a single cycle. Every permutation decomposes uniquely into disjoint cycles, giving $\mathcal{S}=E(\mathcal{C})$ and
$\mathcal{C}(x)=\log\mathcal{S}(x)=\log\frac{1}{1-x}.$
A derangement is a fixed-point-free permutation; the species $\mathrm{Der}\subseteq\mathcal{S}$ satisfies $\mathcal{S}=E\cdot\mathrm{Der}$ since a permutation is a set of fixed points together with a derangement of the rest. Solving for $\mathrm{Der}(x)$ yields
$\mathrm{Der}(x)=\frac{e^{-x}}{1-x}\quad\text{and}\quad|\mathrm{Der}[n]|=n!\sum_{i=0}^{n}\frac{(-1)^{i}}{i!}.$
The species $\mathcal{A}$ of rooted trees plays a key role in this paper. A tree is a connected acyclic simple graph and we let $\mathfrak{a}$ denote the species of trees. A rooted tree is a tree with a distinguished node, called its root; equivalently, a pointed tree, $\mathcal{A}=\mathfrak{a}^{\bullet}$. A rooted tree consists of a root (an $X$-structure) to which is appended a set of rooted trees (an $E(\mathcal{A})$-structure). That is,
$\mathcal{A}\,=\,X\cdot E(\mathcal{A}).$
A slight tweak of the above definition allows us to see rooted trees as a subspecies of functional digraphs. Note that there is exactly one cycle in every connected component of a functional digraph. We identify rooted trees with connected functional digraphs whose only cycle is a loop (a fixed point); this loop plays the role of root, and the edges lose their orientation. Conversely, by putting a loop around the root of a tree and orienting the edges from the leaves towards the root we obtain a connected functional digraph whose only cycle is a loop. A rooted tree and its digraph counterpart are depicted in Figure 2.
3 68915274
3 68915274
Figure 2: The functional digraph of $f=693163933\in\mathrm{End}[9]$, on the left, and the corresponding rooted tree, on the right.
A formula first discovered by Borchardt [3] and later attributed to Cayley [5] shows that there are $n^{n-2}$ labeled trees on $[n]$. Joyal [13] exploited the interplay between rooted trees and endofunctions to prove Cayley’s formula (see also Bergeron et al. [2, Section 2.1]). We recall the core idea, which we extend to two-sort species in Section 6. First, observe that an endofunction is a permutation of rooted trees:
$\mathrm{End}=\mathcal{S}(\mathcal{A}).$ (1)
To see this, let us distinguish two types of points in its functional digraph:
  • Recurrent points: $u$ such that $f^{k}(u)=u$ for some $k\geq 1$; they are the points on cycles. Together they form a set of cycles, that is, a permutation.
  • Nonrecurrent points: $u$ such that $f^{k}(u)\neq u$ for every $k\geq 1$; they are arranged in (rooted) trees each of which hangs from a recurrent point, which can be seen as its root.
The underlying set $U$ is thus partitioned into a set of rooted trees, structured by a permutation of recurrent points. For example, the permutation of the recurrent points of the endofunction in Figure 1 is
$\begin{array}[]{llllll}2\mapsto 8;&3\mapsto 5;&5\mapsto 7;&6\mapsto 6;&7\mapsto 3;&8\mapsto 2.\end{array}$
Joyal proved that doubly-rooted trees and nonempty linear orders of rooted trees are combinatorially equal. Indeed, let $t$ be a structure of $\mathcal{A}^{\bullet}=\mathfrak{a}^{\bullet\bullet}$; that is, a tree where two nodes (not necessarily distinct) are distinguished. Define the spine of $t$ as the directed path from the first distinguished node, the tail, to the second distinguished node, the head. Joyal used the suggestive name “vertebrates” for the species of doubly-rooted trees. The spine determines a nonempty linear order, with a (rooted) tree hanging from each of its nodes. The resulting structure is a nonempty linear order of rooted trees:
$\mathcal{A}^{\bullet}=L_{+}(\mathcal{A}).$
Two species $F$ and $G$ are equipotent, written $F\equiv G$, if $F(x)=G(x)$. Since $L\equiv\mathcal{S}$ we have
$\mathcal{A}^{\bullet}=L_{+}(\mathcal{A})\equiv\mathcal{S}_{+}(\mathcal{A})=\mathrm{End}_{+}$
so $|\mathcal{A}^{\bullet}[n]|=|\mathrm{End}[n]|=n^{n}$ for $n\geq 1$. Finally,
$n^{n}=|\mathcal{A}^{\bullet}[n]|=|\mathfrak{a}^{\bullet\bullet}[n]|=n^{2}\cdot|\mathfrak{a}[n]|$
from which Cayley’s formula immediately follows.
If we assume that $U=\{u_{1},\ldots,u_{n}\}$ is totally ordered, with $u_{1}<u_{2}<\cdots<u_{n}$, then Joyal’s proof induces a bijection between doubly-rooted trees on $U$ and functional digraphs on $U$. In this case, we map the spine $f\in L[U]$ of a doubly-rooted tree to the permutation $g\in\mathcal{S}[U]$ such that $g(u_{i})=u_{j}\Leftrightarrow f(i)=u_{j}$, leaving the rooted trees attached to the spine untouched. This way we obtain a permutation of rooted trees, that is, an endofunction. For instance, the doubly-rooted tree associated with the endofunction of Figure 1 is depicted in Figure 3.
885577663322151511111313441010141499111212
Figure 3: The doubly-rooted tree associated with the endofunction of Figure 1 under Joyal’s bijection. The spine is colored in blue; tail and head are distinguished with one and two circles, respectively.
By adjusting equation (1), one obtains species identities for several classes of endofunctions defined in terms of properties of their functional digraph. For instance, the recurrent points of a connected endofunction form a cyclic permutation, and hence the species of connected endofunctions is $\mathcal{C}(\mathcal{A})$. Similarly, if the recurrent points are structured in a set, we obtain forests (acyclic functional digraphs), $E(\mathcal{A})$. If the permutation of recurrent points is a derangement, we obtain functional digraphs with no fixed points, $\mathrm{Der}(\mathcal{A})$. In some cases, pairing these equations with Cayley’s formula leads to explicit enumeration of the corresponding digraphs. These examples are well known; see Table 1 (A-numbers refer to the OEIS [15]).
Permutations Endofunctions
All structures $\mathcal{S}$ $n!$ $\mathrm{End}$ $n^{n}$
Trees 11 $\delta_{n,0}$ $\mathcal{A}$ $n^{n-1}$
Forests $E$ 11 $E(\mathcal{A})$ $(n+1)^{n-1}$
Connected $\mathcal{C}$ $(n-1)!(1-\delta_{n,0})$ $\mathcal{C}(\mathcal{A})$ A001865
Derangements $\mathrm{Der}$ A000166 $\mathrm{Der}(\mathcal{A})$ $(n-1)^{n}$
Table 1: Enumeration of functional digraphs of permutations and endofunctions, where $\delta$ is the Kronecker delta.
3 Cayley permutations
We now define the species of Cayley permutations [10]. A Cayley permutation on $U$ is a function $f:U\rightarrow[n]$ where $\mathrm{Im}(f)=[k]$ for some $k\leq n=|U|$. The transport of structure along a bijection $\sigma:U\to V$ is given by
$\mathrm{Cay}[\sigma](f)=f\circ\sigma^{-1}.$
Cayley permutations and ballots are combinatorially equal [10]. In the special case where $U=[n]$, we have the inclusions
$\mathcal{S}[n]\subseteq\mathrm{Cay}[n]\subseteq\mathrm{End}[n]$
and all these structures can be identified with their functional digraphs. There is, however, no obvious way to define the functional digraph of Cayley permutations on an arbitrary finite set $U$. Furthermore, the transport of structure on $\mathcal{S}$ and $\mathrm{End}$ is conjugation, which cannot be extended to Cayley permutations due to the requirement $\mathrm{Im}(f)=[k]$. For these reasons, we now assume $\mathbb{L}$-species, defined below.
Let $\mathbb{L}$ denote the category of finite totally ordered sets with order-preserving bijections as morphisms. An $\mathbb{L}$-species is a functor $F:\mathbb{L}\to\mathbb{B}$. Since there is a unique order-preserving bijection between any pair of finite totally ordered sets of the same cardinality, transport of structure is immaterial in the context of $\mathbb{L}$-species. We often assume that the underlying total order is $[n]$ with the standard order and regard permutations and Cayley permutations as subspecies of endofunctions. This allows us to identify all these structures with their functional digraph. Another consequence of having a unique order-preserving bijection between totally ordered sets with the same cardinality is that for $\mathbb{L}$-species $F=G$ if and only if $F(x)=G(x)$. For example, $L=\mathcal{S}$ as $\mathbb{L}$-species even though $L\neq\mathcal{S}$ as $\mathbb{B}$-species. In this case, a permutation $f$ in $\mathcal{S}[n]$ is naturally identified with the linear order $f(1)f(2)\ldots f(n)$ in $L[n]$.
Operations on $\mathbb{B}$-species extend naturally to $\mathbb{L}$-species. New operations also become possible. Given two totally ordered sets $\ell_{1}=(U_{1},\preceq_{1})$ and $\ell_{2}=(U_{2},\preceq_{2})$, their ordinal sum $\ell=\ell_{1}\oplus\ell_{2}$ is the totally ordered set $\ell=(U,\preceq)$ where $U=U_{1}\sqcup U_{2}$, $\preceq$ respects $\preceq_{1}$ and $\preceq_{2}$, and all elements of $U_{1}$ are smaller than the elements of $U_{2}$. The totally ordered sets obtained by adding a new minimum or maximum element to $\ell$ are denoted by $1\oplus\ell$ and $\ell\oplus 1$, respectively. Given a totally ordered set $\ell=(U,\preceq)$, the derivative $F^{\prime}$ of an $\mathbb{L}$-species $F$ is defined by $F^{\prime}[\ell]=F[1\oplus\ell]$. Additionally, the integral of $F$, denoted $\int_{0}^{X}F(T)dT$ or $\int F$, is defined by
[…]
In other words, an $F^{\prime}$-structure is an $F$-structure on a total order endowed with a new minimum, and an $\int\!F$-structure is an $F$-structure on a total order deprived of its minimum. Derivative and integral can be alternatively defined in terms of maximum elements: $F^{\prime}[\ell]=F[\ell\oplus 1]$ and $(\int F)[\ell]=F[\ell\setminus\{\max(\ell)\}]$ for $\ell\neq\emptyset$. Similarly to $\mathbb{B}$-species, operations on $\mathbb{L}$-species behave well with respect to generating series, including a species version of the fundamental theorem of calculus [2, equation 35]:
$\int F^{\prime}=F_{+};\qquad\left(\int F\right)^{\prime}=F;\qquad\left(\int_{0}^{X}F(T)dT\right)(x)=\int_{0}^{x}F(t)dt.$
Let us now return to Cayley permutations. We have observed in Section 2 that functional digraphs of endofunctions are precisely those directed graphs where each node has outdegree one. It is easy to see that a functional digraph is the functional digraph of a permutation if and only if each node has both indegree and outdegree equal to one. The following lemma holds because the internal nodes of $f\in\mathrm{End}[n]$ are labeled by $\mathrm{Im}(f)$, and we omit its proof. It characterizes the functional digraphs of Cayley permutations as those whose internal nodes are labeled with the smallest elements of the underlying total order.
Lemma 3.1.
Let $f\in\mathrm{End}[n]$. Then $f$ is a Cayley permutation with $\mathrm{Im}(f)=[k]$ if and only if the internal nodes in the functional digraph of $f$ have labels $[k]$.
Although $\mathbb{L}$-species address some of the cosmetic issues of Cayley permutations, there are structural problems that require a more sophisticated approach. As an example, consider the following functional digraph:
221133
It is the functional digraph of the endofunction $f$ in $\mathrm{End}[3]$ where $f(1)=f(2)=1$ and $f(3)=3$. Each connected component is a rooted tree that satisfies the requirement of Lemma 3.1: there is only one leaf in the tree on the left, which is labeled with the largest element of the underlying set $\{1,2\}$. However, the resulting functional digraph does not satisfy this property, and indeed $f$ is not a Cayley permutation. More generally, an endofunction is a permutation of rooted trees, but a Cayley permutation is not a permutation of rooted trees of Cayley permutations. Similarly, we cannot use composition to obtain species equations for Cayley permutations whose functional digraph is connected, a forest, or a derangement. Generalizing the equation $\mathcal{S}=E\cdot\mathrm{Der}$ is also not possible: a Cayley permutation is not a pair consisting of the set of its fixed points together with a Cayley permutation without fixed points.
To circumvent these issues, we assume that internal nodes and leaves are nodes of different sorts. This can be formalized using two-sort species.
A two-sort $\mathbb{B}$-species (or $\mathbb{B}$-species of two sorts) is a functor $F:\mathbb{B}\times\mathbb{B}\to\mathbb{B}$, where $\mathbb{B}\times\mathbb{B}$ denotes the product category with pairs of finite sets as objects and pairs of bijections as arrows. Composition of arrows is componentwise and the identity arrow on the pair $(U,V)$ is $\mathrm{id}_{(U,V)}=(\mathrm{id}_{U},\mathrm{id}_{V})$. Two-sort species allow us to build combinatorial structures whose underlying set contains elements of two different sorts: those belonging to $U$ and those belonging to $V$. For a two-sort species $F$, we write $F=F(X,Y)$ to indicate that the two sorts are called $X$ and $Y$. The letters $X$ and $Y$ also serve as singleton species of the corresponding sort:
$X[U,V]$ $=\begin{cases}\{u\},&\text{if $U=\{u\}$ and $V=\emptyset$};\\ \emptyset,&\text{otherwise}.\end{cases}$
$Y[U,V]$ $=\begin{cases}\{v\},&\text{if $U=\emptyset$ and $V=\{v\}$};\\ \emptyset,&\text{otherwise}.\end{cases}$
The generating series of $F(X,Y)$ is
$F(x,y)=\sum_{i,j\geq 0}|F[i,j]|\frac{x^{i}}{i!}\frac{y^{j}}{j!},$
where we let $F[i,j]=F[[i],[j]]$. Most of the operations introduced for unisort species can be naturally extended to the two-sort context, and definitions can be found in the usual references [2, 11]. An example relevant in the next sections is partial differentiation:
$\left(\displaystyle{\frac{\partial{F}}{\partial X}}\right)[U,V]=F[U\cup\{*\},V];\qquad\left(\displaystyle{\frac{\partial{F}}{\partial Y}}\right)[U,V]=F[U,V\cup\{*\}].$
Similarly, we have two pointing operations:
$F^{{}^{\bullet_{X}}}=X\cdot\left(\displaystyle{\frac{\partial{F}}{\partial X}}\right);\qquad F^{{}^{\bullet_{Y}}}=Y\cdot\left(\displaystyle{\frac{\partial{F}}{\partial Y}}\right).$ (2)
Composition of two-sort species is more delicate and requires additional care. We define it below in the only case needed in this paper. Let $F$ be a unisort species and let $G=G(X,Y)$ be a two-sort species. Then $F\circ G=(F\circ G)(X,Y)$ is the two-sort species with structures
$\left(F\circ G\right)[U,V]=\bigcup_{\beta\in\mathrm{Par}[U\sqcup V]}F[\beta]\times\prod_{B\in\beta}G[B\cap U,B\cap V].$
An $\bigl(F\circ G\bigr)$-structure is thus a generalized partition on elements of two sorts, where each block carries a $G$-structure (of two-sorts), and the blocks are structured by the (unisort) species $F$.
4 $R$-recurrent functional digraphs
Recall from Section 2 that rooted trees of one sort are defined by the species equation
$\mathcal{A}(X)\,=\,X\cdot E\bigl(\mathcal{A}(X)\bigr).$
We use the same letter here to denote the two-sort species $\mathcal{A}=\mathcal{A}(X,Y)$ of rooted trees with internal nodes of sort $X$ and leaves of sort $Y$. Then
$\mathcal{A}(X,Y)\,=\,X\cdot E\bigl(\mathcal{A}(X,Y)-X+Y\bigr),$ (3)
where $\mathcal{A}(X,0)=X$ is the tree consisting of a single root (of sort $X$). Indeed—as in the unisort case—a rooted tree of two sorts is obtained by appending a set of rooted trees of two sorts to the root $X$. There is one caveat: when we append a single node $X$, it becomes a leaf in the resulting tree and thus its sort changes from $X$ to $Y$. This explains the additional term X+Y-X+Y in equation (3).
In the unisort case, an endofunction is a permutation of rooted trees, $\mathrm{End}=\mathcal{S}\circ\mathcal{A}$. More precisely, functional digraphs of endofunctions are obtained by structuring the roots of a set of rooted trees in an $\mathcal{S}$-structure. Since this construction preserves internal nodes and leaves, the composition $\mathcal{S}\circ\mathcal{A}(X,Y)$ yields functional digraphs of two sorts where internal nodes have sort $X$ and leaves have sort $Y$. The same holds for the species of Table 1 if we replace $\mathcal{S}$ with $E$ (forests, that is, acyclic functional digraphs), $\mathcal{C}$ (connected functional digraphs), and $\mathrm{Der}$ (derangements, that is, functional digraphs with no fixed points). This brings us to the following definition of $R$-recurrent functional digraphs of two sorts (see also Figure 4).
Definition 4.1.
Let $R$ be a unisort species. The species $\Psi_{R}=\Psi_{R}(X,Y)$ of $R$-recurrent functional digraphs of two sorts is defined by
$\Psi_{R}(X,Y)=R\circ\mathcal{A}(X,Y).$ (4)
In other words, a $\Psi_{R}$-structure is obtained by putting an $R$-structure on the roots of a set of rooted trees of two sorts; in this context, the $R$-structure is called the recurrent part and its points are said to be recurrent. Some significant choices of $R$ are listed in Table 2. Note also that $\Psi_{X}(X,Y)=\mathcal{A}(X,Y)$, which gives
$\Psi_{R}=R\circ\Psi_{X}.$ (5)
1144558899RR1010aa22ddgg6677ff33ccbbee
Figure 4: An $R$-recurrent functional digraph of two sorts. Nodes of sort $X$ (internal nodes) are black and labeled by numbers; nodes of sort $Y$ (leaves) are white and labeled by letters.
The two-sort species $\Psi_{R}$ simultaneously encompasses (subspecies of) Cayley permutations and endofunctions, allowing us to enumerate functional digraphs of both types at once. From an enumerative standpoint, using different sorts for internal nodes and leaves is equivalent to labeling the internal nodes with the smallest elements of the underlying total order (and the remaining elements with the largest), which characterizes functional digraphs of Cayley permutations by Lemma 3.1. On the other hand, using the same sort $X=Y$ for internal nodes and leaves yields all functional digraphs, that is, all endofunctions. We shall make this observation more formal, but first a couple of definitions.
Let $F=F(X,Y)$ be a two-sort species. The unisort species $F(X,X)$ is obtained by identifying the sorts $X$ and $Y$. Note that
$F(X,X)[U]$ $=\bigcup_{U_{1}\sqcup U_{2}=U}F[U_{1},U_{2}]$ (6)
and
$|F(X,X)[n]|$ $=\sum_{i+j=n}\binom{n}{i}|F[i,j]|.$ (7)
Assuming $\mathbb{L}$-species, we also let $\widehat{F}(Z)$ be the unisort species whose structures are
$\widehat{F}[\ell]=\bigcup_{\ell_{1}\oplus\ell_{2}=\ell}F[\ell_{1},\ell_{2}].$ (8)
That is, an $\widehat{F}-$structure on $\ell$ is simply an $F(X,Y)$-structure on $\ell_{1},\ell_{2}]$, where $\ell=\ell_{1}\oplus\ell_{2}$. As a result,
$|\widehat{F}[n]|=\sum_{i+j=n}|F[i,j]|.$ (9)
$R$ $\Psi_{R}(X,Y)$-structures Endofunctions Cayley permutations
$\mathcal{S}$ All structures $\mathrm{End}$ $\mathrm{Cay}$
$X$ Trees $\mathrm{End}_{X}$ $\mathrm{Cay}_{X}$
$E$ Forests $\mathrm{End}_{E}$ $\mathrm{Cay}_{E}$
$\mathcal{C}$ Connected $\mathrm{End}_{\mathcal{C}}$ $\mathrm{Cay}_{\mathcal{C}}$
$\mathrm{Der}\quad$ Derangements $\mathrm{End}_{\mathrm{Der}}$ $\mathrm{Cay}_{\mathrm{Der}}$
Table 2: Unisort species of endofunctions and Cayley permutations for different choices of the recurrent structure $R$ in $\Psi_{R}(X,Y)$.
Now, a $\Psi_{\mathcal{S}}$-structure where internal nodes and leaves have the same sort is simply a functional digraph of one sort, and we have the combinatorial equality
$\Psi_{\mathcal{S}}(X,X)=\mathrm{End}(X).$
On the other hand, by Lemma 3.1 there is a bijection between $\widehat{\Psi}_{\mathcal{S}}[n]$ and $\mathrm{Cay}[n]$. More precisely, structures in $\Psi_{\mathcal{S}}[i,j]$ correspond to functional digraphs of Cayley permutations on $[n]=[i+j]$ that have $i$ internal nodes and $j$ leaves. Therefore,
$\widehat{\Psi}_{\mathcal{S}}(X)=\mathrm{Cay}(X).$
This motivates the following definition.
Definition 4.2.
Let $R$ be a unisort species. The species of $R$-recurrent endofunctions and the species of $R$-recurrent Cayley permutations are defined as follows:
$\mathrm{End}_{R}(X)$ $\,=\,\Psi_{R}(X,X);$
$\mathrm{Cay}_{R}(X)$ $\,=\,\widehat{\Psi}_{R}(X).$
Our main goal is to find a formula for $|\Psi_{R}[i,j]|$. Then, by choosing $R$ as one of $\mathcal{S}$, $X$, $E$, $\mathcal{C}$, or $\mathrm{Der}$, and letting $F=\Psi_{R}$ in equations (7) and (9), we obtain counting formulas for the endofunctions and Cayley permutations of Table 2. Comparing (7) and (9) also shows that the number of endofunctions with $i$ internal nodes and $j$ leaves is simply obtained by multiplying the corresponding number of Cayley permutations by the binomial coefficient $\binom{i+j}{i}$; this additional term accounts for the possibility of labeling internal nodes and leaves of endofunctions without the constraint of using the smallest $i$ labels for the internal nodes.
In Section 3, we observed that neither the equation $\mathrm{End}=\mathcal{S}\circ\mathcal{A}$ nor the equation $\mathcal{S}=E\cdot\mathrm{Der}$ readily extends to Cayley permutations. Both constructions are, however, valid for the two-sort species $\Psi_{\mathcal{S}}$. Indeed, $\Psi_{\mathcal{S}}=\mathcal{S}\circ\Psi_{X}$ by definition, and using $\mathcal{S}=E\cdot\mathrm{Der}$ it follows that
$\Psi_{\mathcal{S}}$ $=(E\cdot\mathrm{Der})\circ\Psi_{X}$
$=E(\Psi_{X})\cdot\mathrm{Der}(\Psi_{X})=\Psi_{E}\cdot\Psi_{\mathrm{Der}}.$
Therefore, in the two-sort sense, functional digraphs are permutations of rooted trees. They are also products of sets of rooted trees (forests) and functional digraphs with no fixed points.
5 A differential equation
Equation (4) defines $\Psi_{R}(X,Y)$ as the composition of $R$ with rooted trees of two sorts. Our goal is a formula for the coefficients $|\Psi_{R}[i,j]|$. While the definition is transparent, it does not readily yield these coefficients. We give an alternative recursive construction leading to a differential equation from which they can be extracted.
[…]
${=}\longleftrightarrow$
${L(X)}\overbrace{}$22774455
[…]
Figure 5: The construction manifested in equation (10).
To build any $\Psi_{R}$-structure, we first build its recurrent part, an $R$-structure of recurrent points of sort $X$. The nonrecurrent points are located on rooted trees of two sorts, each of which hangs from a recurrent point. Such a rooted tree can be obtained recursively. Starting from its root, one repeatedly appends to a specified internal node a path of internal nodes that ends with a leaf; that is, a structure of $L(X)\cdot Y$, which we call a branch. This recursive construction of $\Psi_{R}$-structures is described by the differential equation
${\frac{\partial{}}{\partial Y}}\Psi_{R}(X,Y)=L(X)\cdot\Psi_{R}^{\bullet_{X}}(X,Y),\qquad\Psi_{R}(X,0)=R(X).$ (10)
The initial condition determines the recurrent structure. On the left-hand side of the main equation, we have a structure of $(\partial/\partial Y)\Psi_{R}$; that is, a $\Psi_{R}$-structure with an additional leaf. Equivalently, consider the maximal path starting from the additional leaf (excluded) that consists of internal nodes with indegree at most one. It is an $L(X)$-structure, which is appended to an internal node of a $\Psi_{R}$-structure. By pointing to that internal node we obtain a $\Psi_{R}^{\bullet_{X}}$-structure.
Assuming $\mathbb{L}$-species, we can rewrite equation (10) using the integral operator as
$\Psi_{R}(X,Y)=R(X)+\int L(X)\cdot\Psi_{R}^{\bullet_{X}}(X,Y)dY.$
Figure 5 illustrates this construction. We give a formal proof of equation (10) next. The proof uses the virtual species identity $L^{-1}=1-X$; see [2, Section 2.5] for background on virtual species.
Theorem 5.1.
The two-sort species $\Psi_{R}(X,Y)$ of $R$-recurrent functional digraphs satisfies equation (10).
Proof.
Recall from equations (3) and (5) that $\Psi_{R}=R\circ\Psi_{X}$ and $\Psi_{X}=X\cdot E\bigl(\Psi_{X}-X+Y\bigr)$, where $\Psi_{X}=\mathcal{A}(X,Y)$ and $\Psi_{X}(X,0)=X$. That $\Psi_{R}$ satisfies the initial condition of (10) is easy to see:
$\Psi_{R}(X,0)=R\circ\Psi_{X}(X,0)=R(X).$
Next, we calculate the partial derivatives of $\Psi_{X}$. On differentiating the equation $\Psi_{X}=X\cdot E\bigl(\Psi_{X}-X+Y\bigr)$ we get
${\frac{\partial{\Psi_{X}}}{\partial X}}\,=\,\Bigl(\displaystyle{\frac{\partial{\Psi_{X}}}{\partial X}}+X-1\Bigr)\cdot E(\Psi_{X}-X+Y)$
or, equivalently,
$X\cdot(1-\Psi_{X})\cdot\displaystyle{\frac{\partial{\Psi_{X}}}{\partial X}}\,=\,(1-X)\cdot\Psi_{X}.$
Since $L\cdot(1-X)=1$ and $L_{+}=XL$ we can rewrite this as
$L_{+}(X)\cdot\displaystyle{\frac{\partial{\Psi_{X}}}{\partial X}}=L_{+}\bigl(\Psi_{X}\bigr).$
By a similar calculation we find that
${\frac{\partial{\Psi_{X}}}{\partial Y}}=L_{+}\bigl(\Psi_{X}\bigr)$
and putting these last two equations together we arrive at
${\frac{\partial{\Psi_{X}}}{\partial Y}}=L_{+}(X)\cdot\displaystyle{\frac{\partial{\Psi_{X}}}{\partial X}}.$
Finally, we differentiate $\Psi_{R}=R\circ\Psi_{X}$ to obtain
$\displaystyle{\frac{\partial{\Psi_{R}}}{\partial Y}}$ $=\displaystyle{\frac{\partial{\Psi_{X}}}{\partial Y}}\cdot R^{\prime}(\Psi_{X})$
$=L_{+}(X)\cdot\displaystyle{\frac{\partial{\Psi_{X}}}{\partial X}}\cdot R^{\prime}(\Psi_{X})$
$=L_{+}(X)\cdot\displaystyle{\frac{\partial{\Psi_{R}}}{\partial X}}=L(X)\cdot\Psi_{R}^{\bullet_{X}},$
which concludes the proof. ∎
Equation (10) immediately leads to a recursive formula for $|\Psi_{R}[i,j]|$.
Lemma 5.2.
We have
$|\Psi_{R}[i,j]|=\begin{cases}|R[i]|&j=0;\\ 0&j\geq 1,i=0;\\ i\cdot\bigl(|\Psi_{R}[i,j-1]|+|\Psi_{R}[i-1,j]|\bigr)&i,j\geq 1.\end{cases}$
Proof.
It is easy to see that the first two identities hold. Next, we rewrite equation (10) of Theorem 5.1 as follows:
$\displaystyle{\frac{\partial{\Psi_{R}}}{\partial Y}}=L\cdot X\cdot\displaystyle{\frac{\partial{\Psi_{R}}}{\partial X}}$ $\;\,\iff\;\,\displaystyle{\frac{\partial{\Psi_{R}}}{\partial Y}}\cdot(1-X)=X\displaystyle{\frac{\partial{\Psi_{R}}}{\partial X}}$
$\;\,\iff\;\,\displaystyle{\frac{\partial{\Psi_{R}}}{\partial Y}}=X\cdot\left(\displaystyle{\frac{\partial{\Psi_{R}}}{\partial X}}+\displaystyle{\frac{\partial{\Psi_{R}}}{\partial Y}}\right).$
Identifying coefficients of the last equation gives the third identity. ∎
[…]
This difference of $r$-Stirling numbers counts set partitions of $[n]$ with $m$ blocks such that the numbers $1,2,\ldots,r$ are contained in distinct blocks, but $1,2,\ldots,r,r+1$ are not. In other words, $r$ is maximal such that $1,2,\ldots,r$ appear in distinct blocks. From (11) it follows that
$\genfrac{\{}{\}}{}{}{{n}}{{m}}_{{r}}=m\genfrac{\{}{\}}{}{}{{n-1}}{{m}}_{{r}}+\genfrac{\{}{\}}{}{}{{n-1}}{{m-1}}_{{r}},\quad r<n$ (12)
except $\genfrac{\{}{\}}{}{}{{n}}{{n}}_{{n-1}}=0$, with initial conditions $\genfrac{\{}{\}}{}{}{{n}}{{m}}_{{n}}=\delta_{m,n}$, and $\genfrac{\{}{\}}{}{}{{n}}{{m}}_{{r}}=0$ if $r>n$. In Table 3 we give the numbers $\genfrac{\{}{\}}{}{}{{n}}{{m}}_{{r}}$ for $r=1,2,3,4$, $n=1,2,\ldots,8$ and $m=1,2,\ldots,8$.
$r=1$
1
1 0
1 1 0
1 3 1 0
1 7 6 1 0
1 15 25 10 1 0
1 31 90 65 15 1 0
1 63 301 350 140 21 1 0
$r=2$
0
0 1
0 2 0
0 4 2 0
0 8 10 2 0
0 16 38 18 2 0
0 32 130 110 28 2 0
0 64 422 570 250 40 2 0
$r=3$
0
0 0
0 0 1
0 0 3 0
0 0 9 3 0
0 0 27 21 3 0
0 0 81 111 36 3 0
0 0 243 525 291 54 3 0
$r=4$
0
0 0
0 0 0
0 0 0 1
0 0 0 4 0
0 0 0 16 4 0
0 0 0 64 36 4 0
0 0 0 256 244 60 4 0
Table 3: Triangles of the numbers $\genfrac{\{}{\}}{}{}{{n}}{{k}}_{{r}}$, for $r=1,2,3,4$.
Theorem 5.3.
For $i,j,r\geq 0$, we have
$|\Psi_{R_{r}}[i,j]|=\frac{i!|R[r]|}{r!}\genfrac{\{}{\}}{}{}{{i+j}}{{i}}_{{r}}.$
Proof.
We shall prove that the term on the right-hand side satisfies the recursion of Lemma 5.2 with $R=R_{r}$. The initial conditions are easy to verify, so let us assume that $i,j\geq 1$. We need to prove that
$\frac{i!|R[r]|}{r!}\genfrac{\{}{\}}{}{}{{i+j}}{{i}}_{{r}}=\,i\cdot\left(\frac{(i-1)!|R[r]|}{r!}\genfrac{\{}{\}}{}{}{{i+j-1}}{{i-1}}_{{r}}\;+\;\frac{i!|R[r]|}{r!}\genfrac{\{}{\}}{}{}{{i+j-1}}{{i}}_{{r}}\right).$
But this is equivalent to $\genfrac{\{}{\}}{}{}{{i+j}}{{i}}_{{r}}=\genfrac{\{}{\}}{}{}{{i+j-1}}{{i-1}}_{{r}}+i\genfrac{\{}{\}}{}{}{{i+j-1}}{{i}}_{{r}}$, which follows from (12). ∎
Using Theorem 5.3 and summing over $r$ we arrive at the following corollary.
Corollary 5.4.
For $i,j\geq 0$, we have
$|\Psi_{R}[i,j]|$ $=i!\sum_{r=0}^{i}\frac{|R[r]|}{r!}\genfrac{\{}{\}}{}{}{{i+j}}{{i}}_{{r}}.$ (13)
Recall from Definition 4.2 that $\mathrm{End}_{R}(X)=\Psi_{R}(X,X)$ and $\mathrm{Cay}_{R}(X)=\widehat{\Psi}_{R}(X)$. The number of $R$-recurrent endofunctions and Cayley permutations on $[n]=[i+j]$ are now obtained by combining equation (13) with identities (7) and (9).
Corollary 5.5.
For $n\geq 0$, we have
$|\mathrm{End}_{R}[n]|$ $=\sum_{r=0}^{n}\frac{|R[r]|}{r!}\sum_{i=0}^{n}(n)_{i}\genfrac{\{}{\}}{}{}{{n}}{{i}}_{{r}}$
and
$|\mathrm{Cay}_{R}[n]|$ $=\sum_{r=0}^{n}\frac{|R[r]|}{r!}\sum_{i=0}^{n}i!\genfrac{\{}{\}}{}{}{{n}}{{i}}_{{r}}.$
The expression $(n)_{i}$, above, is the falling factorial; it is defined by
$(n)_{i}=n(n-1)(n-2)\cdots(n-i+1)=\binom{n}{i}i!.$
Letting $R=E_{r}$ in Corollary 5.5, we obtain two identities for the number of forests of $r$ rooted trees of endofunctions and Cayley permutations:
$|\mathrm{End}_{E_{r}}[n]|$ $=\sum_{i=0}^{n}\frac{(n)_{i}}{r!}\genfrac{\{}{\}}{}{}{{n}}{{i}}_{{r}};$
$|\mathrm{Cay}_{E_{r}}[n]|$ $=\sum_{i=0}^{n}\frac{i!}{r!}\genfrac{\{}{\}}{}{}{{n}}{{i}}_{{r}}.$
These allow us to rewrite the formulas of Corollary 5.5 as
$|\mathrm{End}_{R}[n]|$ $=\sum_{r=0}^{n}|R[r]|\cdot|\mathrm{End}_{E_{r}}[n]|;$ (14)
$|\mathrm{Cay}_{R}[n]|$ $=\sum_{r=0}^{n}|R[r]|\cdot|\mathrm{Cay}_{E_{r}}[n]|.$ (15)
The combinatorics underlying equations (14) and (15) is clear: An $\mathrm{End}_{R}$-structure is obtained by putting an $R$-structure over a set of rooted trees. The same holds for $\mathrm{Cay}_{R}$-structures. Furthermore, a forest of endofunctions is a set of rooted trees: $\mathrm{End}_{E}=E\circ\mathrm{End}_{X}$ and $\mathrm{End}_{E_{r}}=E_{r}\circ\mathrm{End}_{X}=\frac{1}{r!}(\mathrm{End}_{X})^{r}$. Thus, hidden underneath equation (14) we find the species composition $\mathrm{End}_{R}=R\circ\mathrm{End}_{X}$, which is is equation (4) when $Y=X$. On the other hand, $\mathrm{Cay}_{E}\neq E\circ\mathrm{Cay}_{X}$.
Alternatively, letting $R=\mathcal{S}_{r}$ we obtain
$|\mathrm{End}_{R}[n]|$ $=\sum_{r=0}^{n}\frac{|R[r]|}{r!}|\mathrm{End}_{S_{r}}[n]|;$
$|\mathrm{Cay}_{R}[n]|$ $=\sum_{r=0}^{n}\frac{|R[r]|}{r!}|\mathrm{Cay}_{S_{r}}[n]|,$
where $|\mathrm{End}_{S_{r}}[n]|$ and $|\mathrm{Cay}_{S_{r}}[n]|$ are simply the number of endofunctions and Cayley permutations with $r$ recurrent points.
By substituting the number $|R[r]|$ of $R$-structures of size $r$ into Theorem 5.3, we obtain counting formulas for the two-sort version of the species of Table 2. Using Corollary 5.5 we also obtain identities for the corresponding unisort structures listed in Table 4. A noteworthy example is a formula for the number of Cayley-derangements, the main result of this paper:
Theorem 5.6.
The number of fixed-point-free Cayley permutations of $[n]$ is
$|\mathrm{Cay}_{\mathrm{Der}}[n]|\,=\,\sum_{r=0}^{n}\frac{|\mathrm{Der}[r]|}{r!}\sum_{i=0}^{n}i!\genfrac{\{}{\}}{}{}{{n}}{{i}}_{{r}}.$
The first few numbers of this sequence are
1,0,1,4,25,184,1617,16492,191721,2503040,36267393,577560596.1,0,1,4,25,184,1617,16492,191721,2503040,36267393,577560596.
At the time of writing, this sequence is not in the OEIS [15].
We highlight three additional noteworthy instances of Corollary 5.5 below.
Proposition 5.7.
The number of Cayley permutations of $[n]$ whose functional digraph is a tree equals the number of all Cayley permutations of [n1][n-1]. That is,
$|\mathrm{Cay}_{X}[n]|=|\mathrm{Cay}[n-1]|.$
The number of Cayley permutations of $[n]$ whose functional digraph is a forest equals
$|\mathrm{Cay}_{E}[n]|$ $=\sum_{r=0}^{n}\frac{1}{r!}\sum_{i=0}^{n}i!\genfrac{\{}{\}}{}{}{{n}}{{i}}_{{r}}.$
The number of Cayley permutations of $[n]$ whose functional digraph is connected equals
$|\mathrm{Cay}_{\mathcal{C}}[n]|$ $=\sum_{r=0}^{n}\frac{1}{r}\sum_{i=0}^{n}i!\genfrac{\{}{\}}{}{}{{n}}{{i}}_{{r}}.$
Proof.
Setting $R=X$ in Corollary 5.5 we get the first identity:
$|\mathrm{Cay}_{X}[n]|=\sum_{r=0}^{n}\frac{|X[r]|}{r!}\sum_{i=0}^{n}i!\genfrac{\{}{\}}{}{}{{n}}{{i}}_{{r}}=\sum_{i=0}^{n}i!\genfrac{\{}{\}}{}{}{{n}}{{i}}_{{1}}=|\mathrm{Cay}[n-1]|,$
where in the last step $\genfrac{\{}{\}}{}{}{{n}}{{i}}_{{1}}$ equals the (n1,i)(n-1,i)th Stirling number of the second kind. The second and third identities are immediate consequences of Corollary 5.5. ∎
In the next section we generalize Joyal’s bijection between doubly-rooted trees and endofunctions to the two-sort setting.
$R$ $|\Psi_{R_{r}}[i,j]|$ $|\mathrm{End}_{R}[n]|$ $|\mathrm{Cay}_{R}[n]|$
All structures $\mathcal{S}$ $i!\genfrac{\{}{\}}{}{}{{i+j}}{{i}}_{{r}}$ $n^{n}$ $|\mathrm{Cay}[n]|$
Trees $X$ $i!\delta_{r,1}\genfrac{\{}{\}}{}{}{{i+j}}{{i}}_{{r}}$ $n^{n-1}$ $|\mathrm{Cay}[n-1]|$
Forests $E$ $\frac{i!}{r!}\genfrac{\{}{\}}{}{}{{i+j}}{{i}}_{{r}}$ $n^{n-2}$ Proposition 5.7
Connected $\mathcal{C}$ $\frac{i!}{r}\genfrac{\{}{\}}{}{}{{i+j}}{{i}}_{{r}}$ A001865 Proposition 5.7
Derangements $\mathrm{Der}$ $|\mathrm{Der}[r]|\frac{i!}{r!}\genfrac{\{}{\}}{}{}{{i+j}}{{i}}_{{r}}$ $(n-1)^{n}$ Theorem 5.6
Table 4: Number of $R$-recurrent functional digraphs with $r$ recurrent points, $i$ internal nodes and $j$ leaves.
6 A two-sort species extension of Joyal’s bijection
Recall from Section 2 Joyal’s bijection between doubly-rooted trees and functional digraphs of endofunctions. It identifies the spine of a doubly-rooted tree with a nonempty permutation, leaving the rooted trees that hang from the spine untouched. This bijection implies the unisort $\mathbb{L}$-species identity
$\mathcal{A}^{\bullet}=\mathrm{End}_{+}.$ (16)
We shall extend Joyal’s construction to the two-sort case by revealing a link between functional digraphs $\Psi_{\mathcal{S}}$ and rooted trees $\Psi_{X}=\mathcal{A}(X,Y)$. In the proof of Proposition 5.1, we showed that
${\frac{\partial{\Psi_{X}}}{\partial Y}}=L_{+}\bigl(\Psi_{X}\bigr).$
Assuming $\mathbb{L}$-species, we have $L=\mathcal{S}$ and hence $L_{+}(\mathcal{A})=\mathcal{S}_{+}(\mathcal{A})=\Psi_{\mathcal{S}_{+}}$. That is,
${\frac{\partial{\Psi_{X}}}{\partial Y}}=\Psi_{\mathcal{S}_{+}}.$ (17)
We have found a combinatorial proof of (17) in the spirit of Joyal’s proof of (16). On the left-hand side, we have a rooted tree of two sorts with an additional leaf. Define its spine as the path from this leaf to the root. Ignoring the initial leaf, this path is a nonempty linear order of (internal) nodes of sort $X$, where a rooted tree of two-sorts hangs from each of its nodes. Identifying the spine with a permutation in the usual way yields a nonempty permutation of rooted trees, that is, a structure of $\mathcal{S}_{+}\circ\Psi_{X}=\Psi_{\mathcal{S}_{+}}$, the species on the right-hand side. In other words, the additional leaf of a structure of $(\partial/\partial Y)\Psi_{X}$ plays the role of tail, while the root plays the role of head. A two-sort vertebrate is born in the process, and by reading a permutation off its spine we obtain a permutation of rooted trees of two sorts; that is, a structure of $\Psi_{\mathcal{S}_{+}}$. This construction—an instance of which can be found in Figure 6—is easily invertible, and thus it yields a combinatorial proof of equation (17).
$y^{\star}$5522443311aabbcceedd
223355ddaabb1144ccee
Figure 6: A rooted tree of two sorts with an additional leaf, denoted $y^{\star}$, and the corresponding functional digraph under the extension of Joyal’s construction. The spine of the tree, in blue, is the linear order 52435243; it is mapped to the recurrent part of the functional digraph on the right.
7 Unisort species of Cayley functional digraphs
In this section, we build on the enumerative results obtained for $\Psi_{R}$ in the two-sort case to obtain species equations for its unisort counterpart $\mathrm{Cay}_{R}$. These rely on the ordinal product of $\mathbb{L}$-species, defined as
$(F\odot G)[\ell]\,=\sum_{\ell=\ell_{1}\oplus\ell_{2}}F[\ell_{1}]\times G[\ell_{2}].$
An $F\odot G$- structure is thus obtained by putting an $F$-structure on the initial segment $\ell_{1}$ and a $G$-structure on the terminal segment $\ell_{2}$.
We shall use $\mathbb{L}$-species to show that
$\mathrm{Cay}_{R}=\frac{1}{2}\sum_{r\geq 0}R_{r}\odot(1+E^{r}\mathrm{Bal}^{r}).$
This specializes to give expressions for the unisort species of Cayley-derangements and the other species in Table 2:
$\mathrm{Cay}_{\mathrm{Der}}=\frac{1}{2}\sum_{r\geq 0}\mathrm{Der}_{r}\odot\left(1+E^{r}\mathrm{Bal}^{r}\right).$ (18)
Let us start with a simple lemma.
Lemma 7.1.
For $r\geq 1$, we have
$\sum_{i\geq 0}i!\cdot\genfrac{\{}{\}}{}{}{{n+r+1}}{{i}}_{{r}}=r!\cdot r\cdot|\bigl(E^{r}\mathrm{Bal}^{r+1}\bigr)[n]|.$
Proof.
The left-hand side counts ballots of $[n+r+1]$ with $i$ blocks, where $r$ is maximal such that $1,\dots,r$ are contained in distinct blocks. Any such ballot is structured
$\beta_{1}\;\{u_{1},\ldots\}\;\beta_{2}\;\{u_{2},\ldots\}\;\dots\;\beta_{r}\;\{u_{r},\ldots\}\;\beta_{r+1},$
where $u_{1}u_{2}\dots u_{r}$ is a permutation of $1,\ldots,r$ and the $\beta_{i}$’s are ballots filling the space between the blocks containing the $u_{i}$’s. The permutation of $1,\ldots,r$ gives the $r!$ term on the right-hand side. Since $r+1$ belongs to one of the blocks containing the $u_{i}$’s, we have $r$ choices to arrange it. Finally, the remaining elements, $r+2,r+3,\ldots,n+r+1$, are arranged in a structure of $E^{r}\cdot\mathrm{Bal}^{r+1}$ (of size $n$): elements in the same block as some $u_{j}$ determine an $E^{r}$-structure, while the other elements determine the ballots $\beta_{1},\dots,\beta_{r+1}$, that is, a structure of $\mathrm{Bal}^{r+1}$. ∎
Theorem 7.2.
We have the $\mathbb{L}$-species identity
$\mathrm{Cay}_{R}=R+\sum_{r\geq 1}(R_{r})^{\bullet}\odot\left(\int E^{r}\mathrm{Bal}^{r+1}\right).$
Proof.
Let $F$ denote the species on the right-hand side. To prove the desired $\mathbb{L}$-species identity, it suffices to show that $|F[n]|=|\mathrm{Cay}_{R}[n]|$ for $n\geq 0$. Expanding the definition of $F$ yields
$|F[n]|$ $=|R[n]|+\sum_{r=1}^{n}|R^{\bullet}[r]|\cdot\left|\left(\int E^{r}\mathrm{Bal}^{r+1}\right)[n-r]\right|$
$=|R[n]|+\sum_{r=1}^{n-1}r|R[r]|\cdot\left|\left(E^{r}\mathrm{Bal}^{r+1}\right)[n-r-1]\right|\cdot\frac{r!}{r!}$
$=|R[n]|+\sum_{r=1}^{n-1}\frac{|R[r]|}{r!}\sum_{i\geq 0}i!\genfrac{\{}{\}}{}{}{{n}}{{i}}_{{r}}$ (by Lemma 7.1)
$=\sum_{r=1}^{n}\frac{|R[r]|}{r!}\sum_{i\geq 0}i!\genfrac{\{}{\}}{}{}{{n}}{{i}}_{{r}}$
$=|\mathrm{Cay}_{R}[n]|$ (by Corollary 5.5)
where in the penultimate equality we used that the term corresponding to $r=n$ in the summation equals
$\frac{|R[n]|}{n!}\sum_{i\geq 0}i!\genfrac{\{}{\}}{}{}{{n}}{{i}}_{{n}}=\frac{n!|R[n]|}{n!}\genfrac{\{}{\}}{}{}{{n}}{{n}}_{{n}}=|R[n]|.$
This concludes the proof. ∎
An arguably simpler species equation for $\mathrm{Cay}_{R}$ is provided in the next result.
Corollary 7.3.
We have
$\mathrm{Cay}_{R}=\frac{1}{2}\sum_{r\geq 0}R_{r}\odot\left(1+E^{r}\mathrm{Bal}^{r}\right).$
Proof.
The $\mathbb{L}$-species identity
$\int E^{r}\mathrm{Bal}^{r+1}=\frac{1}{2r}(E^{r}\mathrm{Bal}^{r}-1).$ (19)
is easily verified using differentiation and the species version of the fundamental theorem of calculus. Next,
$\mathrm{Cay}_{R}$ $=R+\sum_{r\geq 1}(R_{r})^{\bullet}\odot\left(\int E^{r}\mathrm{Bal}^{r+1}\right)$ (by Theorem 7.2)
$=\sum_{r\geq 0}R_{r}+(R_{r})^{\bullet}\odot\left(\int E^{r}\mathrm{Bal}^{r+1}\right)$ (since $R=\sum_{r\geq 0}R_{r}$ & $R^{\bullet}_{0}=\emptyset$)
$=\sum_{r\geq 0}R_{r}+rR_{r}\odot\frac{1}{2r}(E^{r}\mathrm{Bal}^{r}-1)$ (by equation (19))
$=\frac{1}{2}\sum_{r\geq 0}R_{r}\odot\left(1+E^{r}\mathrm{Bal}^{r}\right),$
where the last identity follows by $R_{r}=R_{r}\odot 1$ and linearity of $\odot$. ∎
Finding a combinatorial proof of this corollary remains an open problem.
8 Statistics and asymptotics
Theorem 5.3 provides a closed formula for the number of $\Psi_{R}$-structures with $i$ internal nodes, among which $r$ are recurrent, and $j$ leaves. By suitably summing over $r$, $i$ and $j$, we obtain the distribution of each of these statistics over $R$-recurrent functional digraphs. For instance, the number of $R$-recurrent functional digraphs with $n$ nodes and $r$ recurrent points is
$\sum_{i+j=n}|\Psi_{R_{r}}[i,j]|=\frac{|R[r]|}{r!}\sum_{i=0}^{n}i!\genfrac{\{}{\}}{}{}{{n}}{{i}}_{{r}}.$
The distributions of internal nodes and leaves are obtained similarly. The total number of recurrent points, internal nodes and leaves can be computed in the same fashion by multiplying the term inside the summation by $r$, $i$ or $j$. An example is the total number of recurrent points over $\Psi_{R}$-structures on $[i,j]$:
$\sum_{r=0}^{i}|\Psi_{R_{r}}[i,j]|\cdot r=\sum_{r=0}^{i}\frac{|R[r]|}{r!}i!\genfrac{\{}{\}}{}{}{{i+j}}{{i}}_{{r}}\cdot r.$
Multiplying by $r$ has the same effect as pointing to the $R$-structure. This gives us an elegant species expression for the total number of recurrent points:
$\Psi_{R^{\bullet}}(X,Y),\;\text{ where }|\Psi_{R^{\bullet}}[i,j]|=\sum_{r=0}^{i}|\Psi_{R_{r}}[i,j]|\cdot r.$ (20)
If two species $F$ and $G$ are related by $F=E(G)$, with $G(0)=0$, then $G$ is said to be the species of connected $F$-structures. In this sense, an $F$-structure is a set of $G$-structures each of which is called a connected component. We are by now familiar with the species of connected permutations, that is, cycles $\mathcal{C}$ where $\mathcal{S}=E(\mathcal{C})$. Not every species admits a notion of connected structure in the realm of $\mathbb{B}$-species, but it is always possible [2, Proposition 18], under the assumption $F(0)=1$, to construct a virtual species $G$ such that $F=E(G)$. The species $G$ is uniquely determined and is called the combinatorial logarithm of $F$, denoted by $G=\log F$.
For any unisort species $R$ such that $R(0)=1$, we define the two-sort species of connected $R$-recurrent functional digraphs by letting the recurrent part be equal to the logarithm of $R$:
$\Psi_{\log R}(X,Y).$
Pointing to the $E$-structure in $E(\log R)=R$ has the effect of multiplying by the number of connected components. Therefore, the total number of connected components over $\Psi_{R}$-structures is obtained as $\Psi_{E^{\bullet}(\log R)}$. Note that $E^{\bullet}(X)=X\cdot E(X)$, from which
$E^{\bullet}(\log R)=R\cdot\log R\quad\text{and}\quad\Psi_{E^{\bullet}(\log R)}=\Psi_{R\log R}$ (21)
follows. Finally, Theorem 5.3 gives the total number of connected components over $\Psi_{R}$-structures on $[i,j]$:
$\sum_{r=0}^{i}|\Psi_{R\log R}[i,j]|=\sum_{r=0}^{i}\frac{|(R\log R)[r]|}{r!}i!\genfrac{\{}{\}}{}{}{{i+j}}{{i}}_{{r}}.$ (22)
To illustrate the techniques introduced in this section, we will show that permutations, Cayley permutations and endofunctions share the following property:
The total number of recurrent points over connected structures of a given size equals the total number of structures of that size.
First, we consider permutations. There are (n1)!(n-1)! connected permutations (cycles) of size $n\geq 1$. Since every point in a permutation is recurrent, the total number of recurrent points over connected structures is $(n-1)!\cdot n=n!=|\mathcal{S}[n]|$. To deal with Cayley permutations and endofunctions we use the species of connected functional digraphs, $\Psi_{\mathcal{C}}$. Since $\mathcal{C}^{\prime}=L$, we find that $\mathcal{C}^{\bullet}=X\cdot\mathcal{C}^{\prime}=X\cdot L=L_{+}$. Thus, the total number of recurrent points over $\Psi_{\mathcal{C}}$ is given by
$|\Psi_{\mathcal{C}^{\bullet}}[i,j]|$ $=|\Psi_{L_{+}}[i,j]|=|\Psi_{\mathcal{S}_{+}}[i,j]|.$
Finally, we translate the previous identity to functional digraphs of Cayley permutations and endofunctions to establish the desired property:
$|\mathrm{End}_{\mathcal{C}^{\bullet}}[n]|$ $=|\mathrm{End}_{\mathcal{S}}[n]|=|\mathrm{End}[n]|;$
$|\mathrm{Cay}_{\mathcal{C}^{\bullet}}[n]|$ $=|\mathrm{Cay}_{\mathcal{S}}[n]|=|\mathrm{Cay}[n]|.$
The crucial property that makes the above derivation work is $\mathcal{C}^{\bullet}\equiv\mathcal{S}_{+}$. In other words, the species $F=\mathcal{S}_{+}$ satisfies $(\log F)^{\bullet}\equiv F_{+}$. The solutions to this differential equation are precisely the $k$-colored permutations. That is, $F\equiv\mathcal{S}(kX)$ for some positive integer $k$.
Next, we consider the total number of cycles/connected components over structures in $\Psi_{\mathcal{S}}[i,j]$. It is simply obtained by letting $R=\mathcal{S}$ in equation (21), where $E^{\bullet}(\log\mathcal{S})=E^{\bullet}(\mathcal{C})$ gives the total number of cycles over all permutations. It is well known that the average number of cycles over permutations of $[n]$ is the $n$th harmonic number $H_{n}=\sum_{k=0}^{n}1/k$. To see this, note that $E^{\bullet}(\mathcal{C})=\mathcal{S}\cdot\log(\mathcal{S})=\mathcal{S}\cdot\mathcal{C}$ and
$E^{\bullet}(\mathcal{C})(x)=\bigl(\mathcal{S}\cdot\mathcal{C}\bigr)(x)=\frac{1}{1-x}\cdot\log\left(\frac{1}{1-x}\right).$
Therefore,
$\bigl|E^{\bullet}(\mathcal{C})[n]\bigr|=\bigl|(\mathcal{S}\cdot\mathcal{C})[n]\bigr|=\sum_{k=1}^{n}\binom{n}{k}(n-k)!(k-1)!=n!H_{n}.$
Now, by identity (22), the total number of cycles over $\Psi_{\mathcal{S}}$-structures on $[i,j]$ is
$|\Psi_{\mathcal{S}\log\mathcal{S}}[i,j]|=i!\sum_{r=0}^{i}\frac{r!H_{r}}{r!}\genfrac{\{}{\}}{}{}{{i+j}}{{i}}_{{r}}=i!\sum_{r=0}^{i}H_{r}\genfrac{\{}{\}}{}{}{{i+j}}{{i}}_{{r}}.$
Corollary 5.5 yields the total number of cycles over endofunctions (see A190314 [15]) and Cayley permutations of size $n$:
$|\mathrm{End}_{E^{\bullet}(\mathcal{C})}[n]|$ $=\sum_{i=0}^{n}(n)_{i}\sum_{r=0}^{i}H_{r}\genfrac{\{}{\}}{}{}{{n}}{{i}}_{{r}};$
$|\mathrm{Cay}_{E^{\bullet}(\mathcal{C})}[n]|$ $=\sum_{i=0}^{n}\,i!\,\sum_{r=0}^{i}H_{r}\genfrac{\{}{\}}{}{}{{n}}{{i}}_{{r}}.$
The first few numbers of the latter sequence are
0,1,4,20,126,966,8754,91686,1090578,14528502,214337874,3469418646.0,1,4,20,126,966,8754,91686,1090578,14528502,214337874,3469418646.
At the time of writing, this sequence is not in the OEIS [15]. We conclude this section by presenting two open problems of an asymptotic nature, which we will address in future work.
We opened this paper by recalling the answer to the hat-check problem: the probability that a permutation chosen uniformly at random is a derangement converges to $1/e$ as its size grows to infinity. Numerical evidence computed using Lemma 5.2 suggests that Cayley-derangements have the same asymptotic behavior.
Conjecture 8.1.
We have
$\lim_{n\to\infty}\frac{|\mathrm{Cay}_{\mathrm{Der}}[n]|}{|\mathrm{Cay}[n]|}=\frac{1}{e}.$
The average number of cycles over $\mathrm{End}[n]$ is asymptotic to $\frac{1}{2}(\log(2n)+\gamma)$, where Euler’s constant $\gamma$ is defined as the limiting difference between the harmonic series and the natural logarithm. What is the corresponding limit for the average number of cycles over Cayley permutations? Is there a generalization of Euler’s constant for the average number of cycles over $R$-recurrent functional digraphs?
9 Generalizations
[…]
${\frac{\partial{}}{\partial Y}}\Psi_{R,T}(X,Y)=T(X)\cdot\Psi_{R,T}^{\bullet_{X}}(X,Y),\qquad\Psi_{R,T}(X,0)=R(X).$
For instance, if we let $T=1$ be the empty species, then we obtain structures where all the leaves have distance one from a recurrent point (since each branch $T(X)\cdot Y=Y$ consists of a single leaf). Letting $R=E$, we end up with a “brush-shaped” endofunction, that is, a set of roots (of sort $X$) to each of which is appended a set of leaves (of sort $Y$):
[…]
Note that
$\Psi_{E,1}(X,Y)=E\bigl(X\cdot E(Y)\bigr).$
The corresponding endofunctions of $\mathrm{End}_{E,1}$ and Cayley permutations of $\mathrm{Cay}_{E,1}$ are precisely those $f$ such that $f^{(2)}=f$, where $f^{(2)}=f\circ f$ denotes the composition of $f$ with itself (see also A000248 and A026898 [15]). In other words, $\Psi_{E,1}$ is the species of idempotent functional digraphs of two-sorts.
Let us push this construction a bit further. Setting $R=\mathcal{S}$ has the effect of allowing cycles of any length, yielding maps $f$ where $f^{(k)}=f$ for some $k>1$; in this case,
$\Psi_{\mathcal{S},1}(X,Y)=\mathcal{S}\bigl(X\cdot E(Y)\bigr).$
In the same spirit, we fix $k$ and let
$R=E\circ\sum_{i|k}\mathcal{C}_{i}$
be the species of permutations whose cycle lengths are divisible by $k$. Then $\Psi_{R,1}$ yields maps $f$ such that $f^{(k)}=f$. What other choices of $R$ and $T$ lead to interesting examples?
In Section 4, we have defined the two-sort species $\Psi_{R}=R\circ\mathcal{A}$ whose structures are obtained by arranging a set of rooted trees in an $R$-structure. We can tweak this definition by replacing rooted trees with any other class of trees. For instance, let $\mathcal{B}=\mathcal{B}(X,Y)$ be the two-sort species of rooted trees where each node has at most two (unordered) children. They are obtained as
$\mathcal{B}(X,Y)=X\cdot\Bigl(\bigl(1+X+E_{2}\bigr)\circ\bigl(\mathcal{B}(X,Y)-X+Y\bigr)\bigr).$
Then, $\mathcal{S}\circ\mathcal{B}(X,Y)$ yields functional digraphs where each recurrent point has indegree at most three, and each nonrecurrent point has indegree at most two. In the corresponding endofunctions and Cayley permutations, obtained as usual via (6) and (8), recurrent points have at most three preimages, and nonrecurrent points have at most two preimages (see also A201996 [15]). More generally, rooted trees where each node has at most $k$ children
$\mathcal{K}(X,Y)=X\cdot\Bigl(\bigl(1+X+E_{2}+\cdots+E_{k}\bigr)\circ\bigl(\mathcal{K}(X,Y)-X+Y\bigr)\bigr)$
determine maps where recurrent points have at most $k+1$ preimages and nonrecurrent points have at most $k$ preimages. What can we say about other classes of trees and choices of $R$?